1. 3. Selected year-end data for the Melbourne Company are presented below:              Acid-test ratio 2.5...

1.

3. Selected year-end data for the Melbourne Company are presented below:

            

Acid-test ratio

2.5 to 1

Cost of goods sold

$1,500,000

Current liabilities

$1,800,000

Current ratio

3.0 to 1

The company has no prepaid expenses and inventories remained unchanged during the year. Based on these data, what was the company's inventory turnover ratio for the year?

2. Meryl Company had $540,000 in sales on account last year. The beginning accounts receivable balance was $30,000 and the ending accounts receivable balance was $54,000. What was the company's average collection period (age of receivables)?

3. The Southern Division of Schuler Enterprises recorded operating data as follows for the past year:

Sales

$600,000

Net operating income

75,000

Average operating assets

300,000

Stockholders’ equity

240,000

Residual income

39,000

            What was the margin for the past year?
Answer #1

Part 1

Acid test ratio=(cash & cash equivalents+marketable securities+accounts receivable)/ current liabilities

Therefore assets other than inventory (cash & cash equivalents+marketable securities+accounts receivable) = 2.5 x current liabilities

= 2.5 x 1,800,000 = 4,500,000

Current ratio= Current assets / Current liabilities

Current assets = 3 x current liabilities = 3 x 1,800,000= 5,400,000

Therefore inventory = 5,400,000-4,500,000 = 900,000

Inventory turnover ratio= cost of goods sold/ inventory = 1,500,000/900,000 = 1.667

Part 2

Average collection period= (Average receivables/ Sales) x 365

Average receivables= (30,000+54000)/2 =42000

Average collection period= (42000/ 540000) * 365 = 28 days

Part 3

Margin = Net operating income / sales = (75000/ 600,000) * 100 = 12.5%